two closed vessels A & B of equal volume containing air pressure P 1 and temprature T 1 ARE CONNECTED TO EACH OTHER THROUGH THE NARROW OPEN TUBE. IF THE OF ONE IS NOW MAINTAINED AT T1 AND OTHER AT T2 WHERE T1 > T2 , THEN WHAT WILL BE THE FINAL PRESSURE ?
ans=2p1t2/ti+t2 what is its explanation?

Grade:11

2 Answers

Vikas TU
13 years ago

acc. to ideal gas eqn:

PV=NRT

& it is given that volume is constant means same in both vessels. hence, temp. is directly proprotional to pressure.

as T1>T2 so P1 will also be greater than P2.

that is P1>P2, this is the explanation of your answer..........

 

Vatsal Gheewala
6 years ago
PV=NRT& it is given that volume is constant means same in both vessels. hence, temp. is directly proprotional to pressure.n1 + n2 = n3 + n4 (n1 ,n2 = initial n3 , n4= final )(Po/To) +( Po/To) = (P/To) + (P/T)Now solve this simple equation....wherePo= initial pressureTo= initial temperatureP = final pressureT = final temperature